# RedactOR: An LLM-Powered Framework for Automatic Clinical Data De-Identification

### v1

@@ Abstract
Ensuring clinical data privacy while preserving utility is critical for AI-driven healthcare and data analytics. Existing de-identification (De-ID) methods, including rule-based techniques, deep learning models, and large language models (LLMs), often suffer from recall errors, limited generalization, and inefficiencies, limiting their real-world applicability. We propose a fully automated, multi-modal framework, RedactOR for de-identifying structured and unstructured electronic health records, including clinical audio records. Our framework employs cost-efficient De-ID strategies, including intelligent routing, hybrid rule and LLM based approaches, and a two-step audio redaction approach. We present a retrieval-based entity relexicalization approach to ensure consistent substitutions of protected entities, thereby enhancing data coherence for downstream applications. We discuss key design desiderata, de-identification and relexicalization methodology, and modular architecture of RedactOR and its integration with the Oracle Health Clinical AI system. Evaluated on the i2b2 2014 De-ID dataset using standard metrics with strict recall, our approach achieves competitive performance while optimizing token usage to reduce LLM costs. Finally, we discuss key lessons and insights from deployment in real-world AI-driven healthcare data pipelines.
RedactOR: An LLM-Powered Framework for Automatic Clinical Data De-Identification
@@ 1 Introduction
We propose RedactOR, a fully automated, multi-modal framework for de-identifying structured and unstructured patient records, including clinical audio records. RedactOR combines LLM-based processing for unstructured text with rule-based handling of structured data to achieve low cost and latency, and extends text de-identification for audio redaction. Our framework includes a novel retrieval-based entity relexicalization component to ensure consistent PHI/PII replacement, enhancing coherence and privacy. We present key design requirements (§3.1), de-identification and relexicalization methodology and architecture (§3), and integration with the Oracle Health Clinical AI system (§A.8). We demonstrate that our framework outperforms other LLM-based approaches and achieves performance comparable to specialized, closed-source solutions while remaining adaptable through prompt engineering – eliminating reliance on large annotated datasets (§4), and highlight lessons and insights from 12+ months of deployment in real-world clinical AI system (§5).
Figure 1: Architectural overview of RedactOR
@@ 3.1 Design Desiderata
RedactOR is designed around three core principles: scalability, adaptability, and cost-efficiency. Scalability is achieved through an end-to-end automation pipeline, enabling efficient processing of structured and unstructured clinical data while minimizing computational overhead via intelligent routing and LLM-based De-ID strategies. Adaptability is ensured through a schema-agnostic processing architecture, facilitating seamless integration across heterogeneous EHR formats and multimodal data sources (text and audio) without the need for dataset-specific fine-tuning. Cost-efficiency is realized through token usage optimization in text-based De-ID and leveraging text-based entity extraction for audio, thereby eliminating the need for computationally expensive, audio-specific de-identification models. Additionally, retrieval-based re-lexicalization enhances contextual consistency in PHI/PII replacements, preserving both privacy and downstream utility, making the system highly effective for real-world AI-driven healthcare applications.
@@ 3.2 Architecture Overview
RedactOR consists of three main components: (i) Auto De-ID, Audio De-ID, and Auto Relexicalizer (see Figure 1).
@@
By integrating Auto De-ID for unstructured text, Auto Relexicalizer for realistic entity replacement, and Audio De-ID for speech data, RedactOR provides a scalable, adaptable, and cost-effective De-ID pipeline that secures both text and audio data while preserving its utility.
@@ 4 Experiments
We present the results of evaluating RedactOR against other LLM-based approaches and specialized, closed-source commercial solutions over a publicly available medical record dataset. For parity with other methods, we turn off the Auto Relexicalizer component. We set chunk size ($\omega$) to 256 and number of passes ($p$) to 2.
@@ 4.2 Comparative De-ID methods
We evaluated RedactOR against recent LLM-based methods, Yashwanth and Shettar (2024) (with their two prompt variants ‘brief’ and ‘detailed’) and Altalla’ et al. (2025), as well as commercial De-ID APIs from AWS (Amazon Web Services) AWS (2025) and JSL (John Snow Labs) Kocaman et al. (2023, 2025). For a fair comparison, we used GPT-4o OpenAI (2025) for all LLM-based methods. See §A.3 for additional details.
@@ 4.3 Evaluation Methods
Entity Type RedactOR AWS JSL
@@ 4.4 Comparison with LLM-Based De-ID Frameworks
As shown in Table 1, a comparison with other LLM-based methods Yashwanth and Shettar (2024); Altalla’ et al. (2025) indicates that RedactOR outperforms existing methods, achieving the highest F1-score of 0.9646 with a well-balanced precision and recall. The high recall can be attributed to RedactOR’s multi-chunk and multi-pass strategy, which systematically refines entity detection by iteratively masking extracted entities and forcing the model to focus on overlooked PHI. Similarly, in terms of precision, RedactOR outperforms Yashwanth and Shettar (2024), highlighting the effectiveness of its context-aware entity extraction. By leveraging contextual clues and maintaining intra-document consistency, RedactOR reduces false positives, whereas single-pass prompting methods tend to over-redact ambiguous terms.
While Yashwanth and Shettar (2024)’s Detailed version achieves higher recall than the Brief, it does so at the expense of precision. This highlights a fundamental trade-off in zero-shot protected terms extraction: models optimized for recall often over-mask non-protected terms, leading to reduced utility of the redacted text. RedactOR strikes a balance between recall and precision, making it more suitable for real-world clinical applications where both PHI/PII removal and utility are essential.
@@ 4.5 Comparisons with Commercial APIs
Unlike prior zero-shot LLM-based approaches, commercial De-ID APIs (e.g., AWS, JSL) are fine-tuned on proprietary clinical datasets. Table 1 shows that while RedactOR does not surpass JSL in recall, it achieves higher precision and a comparable F1-score. This suggests that while JSL benefits from domain-specific fine-tuning, RedactOR’s context-aware extraction minimizes false positives, leading to more reliable entity masking.
Entity Type RedactOR AWS JSL
@@
Table 3: All-or-nothing recalls constrained on entity types for RedactOR and commercial models.
Table 2 presents a breakdown of performance by entity types. RedactOR demonstrates high precision and strong recall across all entities, particularly excelling on CONTACT and PERSON entities. RedactOR achieves perfect recall on CONTACT entities, outperforming AWS and JSL, and shows competitive performance on PERSON and DATE entities. However, it underperforms on LOCATION and ID, presumably due to the structural variability of PHI in clinical texts. LOCATION entities, in particular, often appear within complex sentence structures, posing challenges to generic LLM-based masking. This suggests the need for instruction updates for these entities or a specialized LLM.
@@ 4.6 All-or-Nothing Recall Results
Table 3 shows performance by PHI/PII entity type with all-or-nothing recall, a stricter metric requiring both correct entity type and position alignment. These results highlight RedactOR’s strength in high-precision redaction for certain entities while emphasizing the advantage of domain-tuned models for broader recall coverage.
RedactOR outperforms other methods using the same underlying LLM (Table 1) and approaches the performance of specialized models like JSL. Specifically, it excels in CONTACT and ID compared to commercial methods. RedactOR’s multi-pass extraction and context-aware masking enhance the LLM’s effectiveness, showcasing its strength without specialized fine-tuning. However, JSL leads in most other entity types, achieving the highest overall recall. The recall gap – especially for complex entities like DATE and LOCATION – highlights the need for improved prompt instructions and possibly a specialized LLM for de-identification to match domain-specific systems.
@@ 4.7 Ablation Study with Open-Source Model
To demonstrate the adaptability of RedactOR to open-source models and evaluate the benefit of its multi-pass de-identification strategy, we conducted an ablation study using LLaMA-3.2-3B-Instruct MetaAI (2024) – a compact, publicly available LLM.
@@
These trends suggest that the number of passes is a critical, model-dependent hyperparameter: lightweight models like LLaMA-3.2-3B benefit from 2–3 passes, while larger models may reach optimal performance sooner. RedactOR supports this flexibility by treating the pass count as a configurable parameter, allowing practitioners to trade off between computational cost and de-identification completeness depending on the capacity of the underlying LLM.
@@ 5 Deployment Lessons and Insights
In the course of developing and deploying RedactOR, we realized that efficient scaling of de-identification is crucial to handle the large volume of healthcare data post-deployment, including audio files, SOAP notes, and longitudinal records with thousands of FHIR resources Bender and Sartipi (2013). Once the service is launched, a continuous influx of data follows, and as the products expand, the ability to optimize processing at scale becomes critical. Simple yet powerful reductions in computation and processing can significantly impact efficiency, cost, and system performance.
One key optimization we employed was reducing token usage. RedactOR extracts only PHI/PII entities and their position hints, minimizing the number of LLM’s output tokens by approximately 50%. This reduction not only lowers processing costs but also decreases latency, ensuring that De-ID remains accurate and efficient at scale.
@@ 6 Conclusion And Future Directions
Motivated by the need for protecting patient privacy while enabling utility, we presented RedactOR  a multi-modal, scalable, flexible, and cost-efficient LLM-powered framework for clinical data de-identification, and demonstrated its efficacy in de-identifying 33 PHI/PII entities over the i2b2 dataset. We showed that our approach outperforms other LLM-based methods and achieves performance comparable to specialized, closed-source solutions. Further, RedactOR supports relexicalizing redacted entities with contextually consistent alternatives, enhancing data usability and strengthening privacy. By presenting the methodology, technical architecture, and lessons learned from over 12 months of production deployment, we hope that the insights and experience from our work are useful for researchers and practitioners working on clinical AI systems.
There are several avenues for future work. The variability in healthcare datasets across institutions affects generalizability, necessitating adaptive prompting techniques. While our method excels in detecting ID, PERSON, and DATE entities, it may require further refinement of entity-specific LLM instructions (e.g., for address- and occupation-related entities). More broadly, RedactOR’s generalizability can be enhanced across diverse institutional datasets by integrating domain-adaptive prompt enhancements.
@@ Acknowledgments
We would like to thank other members of Oracle Health AI for their collaboration while deploying RedactOR in production, and Brent Beardsley, Michelle Chen, Raefer Gabriel, Kent Grueneich, Neil Hauge, Gyan Shankar, Brad Jacobs, Shirley Liu, Cody Maheu, Virendra Marathe, Vishal Vishnoi and Mark Johnson for insightful feedback and discussions.
@@ A.1 Ethics Statement
Given the sensitivity of De-ID data pipelines, we do not release the prompt verbatim or source code of RedactOR to prevent potential privacy risks and attacks. However, to support transparency and reproducibility, we provide descriptions of individual components, including retrieval-based relexicalization, hybrid rule/LLM logic, intelligent routing, and a two-step audio redaction process, in the pipeline. This approach enables secure replication of our methodology while safeguarding patient confidentiality in our system.
@@
The de-identified data produced by RedactOR is intended for a variety of critical use cases within AI-driven healthcare systems. First, it serves as a valuable resource for debugging production issues, enabling engineers and data scientists to analyze system behavior and identify root causes of errors without compromising patient privacy. Second, the data supports understanding production model behavior, providing insights into model performance, biases, and failure modes, which guide iterative improvements and model refinements. Finally, the de-identified dataset can be leveraged for training and evaluating downstream machine learning models, including clinical documentation automation, clinical named entity recognition, and “needle in a haystack” tasks such as rare condition detection or retrieval of highly specific information from longitudinal records. Additionally, de-identified data is essential for conducting R&D and facilitates collaboration with external researchers and clinicians, enabling innovation while ensuring compliance with privacy regulations. These applications illustrate the dual importance of ensuring privacy while maintaining data utility for real-world healthcare advancements.
@@ A.3 All Models
    RedactOR (ours): The Auto De-ID model’s outputs are post-processed to align with our predefined PHI/PII entity categories. We use the following configuration parameters to ensure consistency across experiments: max_passes = 2, max_words = 256, and temperature = 0.
@@ A.7 Proposed Metrics for Relexicalization
Figure 5: Diagram explaining the integration of our RedactOR framework with Oracle Health Clinical AI System
A.8 Integration of the RedactOR with Oracle Health & Clinical AI System
Our RedactOR framework is integrated into Oracle Health Clinical AI system to facilitate the privacy-preserving processing of longitudinal EHRs, ambient intelligence data, and conversational AI outputs. The system operates autonomously, as the Production Environment is inaccessible to any user group, ensuring a fully automated pipeline. A dedicated worker module, running continuously, monitors the Datasink, retrieves unprocessed files, submits them to the RedactOR service for PHI/PII removal, and securely transfers the de-identified records to the Data Science Lab Environment. A one-way policy enforces strict data flow control, guaranteeing that only de-identified data is accessible within a secure research environment, where authorized users interact with it via an SSH-secured virtual machine, preserving both data integrity and analytical utility.
A.9 End to End Example Illustrating Text De-identification and Relexicalization using RedactOR

### v2

# RedactX: An LLM-Powered Framework for Automatic Clinical Data De-Identification

@@ Abstract
Ensuring clinical data privacy while preserving utility is critical for AI-driven healthcare and data analytics. Existing de-identification (De-ID) methods, including rule-based techniques, deep learning models, and large language models (LLMs), often suffer from recall errors, limited generalization, and inefficiencies, limiting their real-world applicability. We propose a fully automated, multi-modal framework, RedactX for de-identifying structured and unstructured electronic health records, including clinical audio records. Our framework employs cost-efficient De-ID strategies, including intelligent routing, hybrid rule and LLM based approaches, and a two-step audio redaction approach. We present a retrieval-based entity relexicalization approach to ensure consistent substitutions of protected entities, thereby enhancing data coherence for downstream applications. We discuss key design desiderata, de-identification and relexicalization methodology, and modular architecture of RedactX and its integration with the Oracle Health Clinical AI system. Evaluated on the i2b2 2014 De-ID dataset using standard metrics with strict recall, our approach achieves competitive performance while optimizing token usage to reduce LLM costs. Finally, we discuss key lessons and insights from deployment in real-world AI-driven healthcare data pipelines.
RedactX: An LLM-Powered Framework for Automatic Clinical Data De-Identification
@@ 1 Introduction
We propose RedactX, a fully automated, multi-modal framework for de-identifying structured and unstructured patient records, including clinical audio records. RedactX combines LLM-based processing for unstructured text with rule-based handling of structured data to achieve low cost and latency, and extends text de-identification for audio redaction. Our framework includes a novel retrieval-based entity relexicalization component to ensure consistent PHI/PII replacement, enhancing coherence and privacy. We present key design requirements (§3.1), de-identification and relexicalization methodology and architecture (§3), and integration with the Oracle Health Clinical AI system (§A.8). We demonstrate that our framework outperforms other LLM-based approaches and achieves performance comparable to specialized, closed-source solutions while remaining adaptable through prompt engineering – eliminating reliance on large annotated datasets (§4), and highlight lessons and insights from 12+ months of deployment in real-world clinical AI system (§5).
Figure 1: Architectural overview of RedactX
@@ 3.1 Design Desiderata
RedactX is designed around three core principles: scalability, adaptability, and cost-efficiency. Scalability is achieved through an end-to-end automation pipeline, enabling efficient processing of structured and unstructured clinical data while minimizing computational overhead via intelligent routing and LLM-based De-ID strategies. Adaptability is ensured through a schema-agnostic processing architecture, facilitating seamless integration across heterogeneous EHR formats and multimodal data sources (text and audio) without the need for dataset-specific fine-tuning. Cost-efficiency is realized through token usage optimization in text-based De-ID and leveraging text-based entity extraction for audio, thereby eliminating the need for computationally expensive, audio-specific de-identification models. Additionally, retrieval-based re-lexicalization enhances contextual consistency in PHI/PII replacements, preserving both privacy and downstream utility, making the system highly effective for real-world AI-driven healthcare applications.
@@ 3.2 Architecture Overview
RedactX consists of three main components: (i) Auto De-ID, Audio De-ID, and Auto Relexicalizer (see Figure 1).
@@
By integrating Auto De-ID for unstructured text, Auto Relexicalizer for realistic entity replacement, and Audio De-ID for speech data, RedactX provides a scalable, adaptable, and cost-effective De-ID pipeline that secures both text and audio data while preserving its utility.
@@ 4 Experiments
We present the results of evaluating RedactX against other LLM-based approaches and specialized, closed-source commercial solutions over a publicly available medical record dataset. For parity with other methods, we turn off the Auto Relexicalizer component. We set chunk size ($\omega$) to 256 and number of passes ($p$) to 2.
@@ 4.2 Comparative De-ID methods
We evaluated RedactX against recent LLM-based methods, Yashwanth and Shettar (2024) (with their two prompt variants ‘brief’ and ‘detailed’) and Altalla’ et al. (2025), as well as commercial De-ID APIs from AWS (Amazon Web Services) AWS (2025) and JSL (John Snow Labs) Kocaman et al. (2023, 2025). For a fair comparison, we used GPT-4o OpenAI (2025) for all LLM-based methods. See §A.3 for additional details.
@@ 4.3 Evaluation Methods
Entity Type RedactX AWS JSL
@@ 4.4 Comparison with LLM-Based De-ID Frameworks
As shown in Table 1, a comparison with other LLM-based methods Yashwanth and Shettar (2024); Altalla’ et al. (2025) indicates that RedactX outperforms existing methods, achieving the highest F1-score of 0.9646 with a well-balanced precision and recall. The high recall can be attributed to RedactX’s multi-chunk and multi-pass strategy, which systematically refines entity detection by iteratively masking extracted entities and forcing the model to focus on overlooked PHI. Similarly, in terms of precision, RedactX outperforms Yashwanth and Shettar (2024), highlighting the effectiveness of its context-aware entity extraction. By leveraging contextual clues and maintaining intra-document consistency, RedactX reduces false positives, whereas single-pass prompting methods tend to over-redact ambiguous terms.
While Yashwanth and Shettar (2024)’s Detailed version achieves higher recall than the Brief, it does so at the expense of precision. This highlights a fundamental trade-off in zero-shot protected terms extraction: models optimized for recall often over-mask non-protected terms, leading to reduced utility of the redacted text. RedactX strikes a balance between recall and precision, making it more suitable for real-world clinical applications where both PHI/PII removal and utility are essential.
@@ 4.5 Comparisons with Commercial APIs
Unlike prior zero-shot LLM-based approaches, commercial De-ID APIs (e.g., AWS, JSL) are fine-tuned on proprietary clinical datasets. Table 1 shows that while RedactX does not surpass JSL in recall, it achieves higher precision and a comparable F1-score. This suggests that while JSL benefits from domain-specific fine-tuning, RedactX’s context-aware extraction minimizes false positives, leading to more reliable entity masking.
Entity Type RedactX AWS JSL
@@
Table 3: All-or-nothing recalls constrained on entity types for RedactX and commercial models.
Table 2 presents a breakdown of performance by entity types. RedactX demonstrates high precision and strong recall across all entities, particularly excelling on CONTACT and PERSON entities. RedactX achieves perfect recall on CONTACT entities, outperforming AWS and JSL, and shows competitive performance on PERSON and DATE entities. However, it underperforms on LOCATION and ID, presumably due to the structural variability of PHI in clinical texts. LOCATION entities, in particular, often appear within complex sentence structures, posing challenges to generic LLM-based masking. This suggests the need for instruction updates for these entities or a specialized LLM.
@@ 4.6 All-or-Nothing Recall Results
Table 3 shows performance by PHI/PII entity type with all-or-nothing recall, a stricter metric requiring both correct entity type and position alignment. These results highlight RedactX’s strength in high-precision redaction for certain entities while emphasizing the advantage of domain-tuned models for broader recall coverage.
RedactX outperforms other methods using the same underlying LLM (Table 1) and approaches the performance of specialized models like JSL. Specifically, it excels in CONTACT and ID compared to commercial methods. RedactX’s multi-pass extraction and context-aware masking enhance the LLM’s effectiveness, showcasing its strength without specialized fine-tuning. However, JSL leads in most other entity types, achieving the highest overall recall. The recall gap – especially for complex entities like DATE and LOCATION – highlights the need for improved prompt instructions and possibly a specialized LLM for de-identification to match domain-specific systems.
@@ 4.7 Ablation Study with Open-Source Model
To demonstrate the adaptability of RedactX to open-source models and evaluate the benefit of its multi-pass de-identification strategy, we conducted an ablation study using LLaMA-3.2-3B-Instruct MetaAI (2024) – a compact, publicly available LLM.
@@
These trends suggest that the number of passes is a critical, model-dependent hyperparameter: lightweight models like LLaMA-3.2-3B benefit from 2–3 passes, while larger models may reach optimal performance sooner. RedactX supports this flexibility by treating the pass count as a configurable parameter, allowing practitioners to trade off between computational cost and de-identification completeness depending on the capacity of the underlying LLM.
@@ 5 Deployment Lessons and Insights
In the course of developing and deploying RedactX, we realized that efficient scaling of de-identification is crucial to handle the large volume of healthcare data post-deployment, including audio files, SOAP notes, and longitudinal records with thousands of FHIR resources Bender and Sartipi (2013). Once the service is launched, a continuous influx of data follows, and as the products expand, the ability to optimize processing at scale becomes critical. Simple yet powerful reductions in computation and processing can significantly impact efficiency, cost, and system performance.
One key optimization we employed was reducing token usage. RedactX extracts only PHI/PII entities and their position hints, minimizing the number of LLM’s output tokens by approximately 50%. This reduction not only lowers processing costs but also decreases latency, ensuring that De-ID remains accurate and efficient at scale.
@@ 6 Conclusion And Future Directions
Motivated by the need for protecting patient privacy while enabling utility, we presented RedactX  a multi-modal, scalable, flexible, and cost-efficient LLM-powered framework for clinical data de-identification, and demonstrated its efficacy in de-identifying 33 PHI/PII entities over the i2b2 dataset. We showed that our approach outperforms other LLM-based methods and achieves performance comparable to specialized, closed-source solutions. Further, RedactX supports relexicalizing redacted entities with contextually consistent alternatives, enhancing data usability and strengthening privacy. By presenting the methodology, technical architecture, and lessons learned from over 12 months of production deployment, we hope that the insights and experience from our work are useful for researchers and practitioners working on clinical AI systems.
There are several avenues for future work. The variability in healthcare datasets across institutions affects generalizability, necessitating adaptive prompting techniques. While our method excels in detecting ID, PERSON, and DATE entities, it may require further refinement of entity-specific LLM instructions (e.g., for address- and occupation-related entities). More broadly, RedactX’s generalizability can be enhanced across diverse institutional datasets by integrating domain-adaptive prompt enhancements.
@@ Acknowledgments
We would like to thank other members of Oracle Health AI for their collaboration while deploying RedactX in production, and Weifeng Bao, Brent Beardsley, Michelle Chen, Dipankar Das, Long Duong, Raefer Gabriel, Kent Grueneich, Neil Hauge, Bhagya Hettige, Brad Jacobs, Mark Johnson, Shirley Liu, Cody Maheu, Atri Mandal, Virendra Marathe, Kiran Rama, Amitabh Saikia, Tushar Shandhilya, Gyan Shankar, and Vishal Vishnoi for insightful feedback and discussions.
@@ A.1 Ethics Statement
Given the sensitivity of De-ID data pipelines, we do not release the prompt verbatim or source code of RedactX to prevent potential privacy risks and attacks. However, to support transparency and reproducibility, we provide descriptions of individual components, including retrieval-based relexicalization, hybrid rule/LLM logic, intelligent routing, and a two-step audio redaction process, in the pipeline. This approach enables secure replication of our methodology while safeguarding patient confidentiality in our system.
@@
The de-identified data produced by RedactX is intended for a variety of critical use cases within AI-driven healthcare systems. First, it serves as a valuable resource for debugging production issues, enabling engineers and data scientists to analyze system behavior and identify root causes of errors without compromising patient privacy. Second, the data supports understanding production model behavior, providing insights into model performance, biases, and failure modes, which guide iterative improvements and model refinements. Finally, the de-identified dataset can be leveraged for training and evaluating downstream machine learning models, including clinical documentation automation, clinical named entity recognition, and “needle in a haystack” tasks such as rare condition detection or retrieval of highly specific information from longitudinal records. Additionally, de-identified data is essential for conducting R&D and facilitates collaboration with external researchers and clinicians, enabling innovation while ensuring compliance with privacy regulations. These applications illustrate the dual importance of ensuring privacy while maintaining data utility for real-world healthcare advancements.
@@ A.3 All Models
    RedactX (ours): The Auto De-ID model’s outputs are post-processed to align with our predefined PHI/PII entity categories. We use the following configuration parameters to ensure consistency across experiments: max_passes = 2, max_words = 256, and temperature = 0.
@@ A.7 Proposed Metrics for Relexicalization
Figure 5: Diagram explaining the integration of our RedactX framework with Oracle Health Clinical AI System
A.8 Integration of the RedactX with Oracle Health & Clinical AI System
Our RedactX framework is integrated into Oracle Health Clinical AI system to facilitate the privacy-preserving processing of longitudinal EHRs, ambient intelligence data, and conversational AI outputs. The system operates autonomously, as the Production Environment is inaccessible to any user group, ensuring a fully automated pipeline. A dedicated worker module, running continuously, monitors the Datasink, retrieves unprocessed files, submits them to the RedactX service for PHI/PII removal, and securely transfers the de-identified records to the Data Science Lab Environment. A one-way policy enforces strict data flow control, guaranteeing that only de-identified data is accessible within a secure research environment, where authorized users interact with it via an SSH-secured virtual machine, preserving both data integrity and analytical utility.
A.9 End to End Example Illustrating Text De-identification and Relexicalization using RedactX